OpenStudy (anonymous):

find the exact value of each of the remaining trigonometric functions of theta sec(theta)=-8, sin(theta)>0

OpenStudy (anonymous):

magic

OpenStudy (anonymous):

\[\sec(\theta) = -8\]

OpenStudy (anonymous):

recall that \[\sec(\theta) = \frac {1}{\cos(\theta)}\]

OpenStudy (anonymous):

so, \[\frac {1}{\cos(\theta)} = -8 \rightarrow \frac {1}{-8} = \cos(\theta) \rightarrow \cos^{-1}(\frac{1}{-8}) = \theta\]